OpenStudy (anonymous):

how many kilograms of water are in a 60-kg person?

OpenStudy (anonymous):

There's 36 kg of water in a 50 kg person so W = k*B. Solve for "k" using "A person who weighs 50 kg contains 36kg of water. 36 = k*50 k = 36/50 Equation: W = (36/50)B W = (36/50)60 W = 43.2kg *thanks to stan h*